OpenStudy (anonymous):

Find the slope and the y intercept f(x)=-10-3 the slope is? the y intercept is?

jimthompson5910 (jim_thompson5910):

Is it supposed to be f(x) = -10x-3 ?

jimthompson5910 (jim_thompson5910):

If so, then.... f(x) = -10x-3 is in the form f(x) = mx+b which is basically the same as y = mx+b The general equation y = mx+b has a slope of m and a y intercept of (0, b) In the case of f(x) = -10x-3, we can see that m = -10 and b = -3 So the slope is -10 and the y intercept is (0, -3)

OpenStudy (anonymous):

Your a math whiz, you did that rather fast.You earned your medal.Thanks!

jimthompson5910 (jim_thompson5910):

yw